OpenStudy (anonymous):

solve the system of equations , out in ordered pair .. 6x+5y=61 -5x+5y=-5

OpenStudy (anonymous):

*Put

jimthompson5910 (jim_thompson5910):

Multiply everything in the second equation by -1 to get 5x - 5y = 5 So we now have the system 6x+5y=61 5x - 5y = 5 which adds to 6x+5y=61 5x - 5y = 5 ----------- 11x+0y = 66 or just 11x = 66 See where to go from here?

OpenStudy (anonymous):

Put it in ordered pairs .. so X=6 how do i solve for y ?

jimthompson5910 (jim_thompson5910):

If x = 6, then... 6x+5y=61 6*6+5y=61 36 + 5y = 61 Now solve for y 36 + 5y = 61 5y = 61-36 5y = 25 y = 25/5 y = 5 Note: you can use either equation to find y